OpenStudy (anonymous):

Consider this equation: 4x – 5 + 6x = 6 Which is an equivalent equation after combining like terms?

OpenStudy (anonymous):

\(\Large 4x – 5 + 6x = 6\) \(\Large 10x-5-6=0\) \(\Large 10x-11=0\)